A 0.151kg glider is moving to the right on a frictionless, horizontal air track with a speed of 0.840m/s . It has a head-on collision with a 0.302kg glider that is moving to the left with a speed of 2.26m/s . Suppose the collision is elastic.

Part A

Find the magnitude of the final velocity of the 0.151kg glider.

Part B

Find the magnitude of the final velocity of the 0.302kg glider
Answer #1

Here , let the speed of 0.151 kg is v1

the final speed of 0.302 kg is v2

Using conservation of momentum

0.151 *v1 + 0.302 * v2 = 0.151 *0.84 - 2.26 * 0.302

0.151 *v1 + 0.302 * v2 = -0.56     ---(1)

as the collision is elastic

v2 - v1 = (2.26 + 0.84)   ----(2)

solving 1 and 2

v1 = -3.3 m/s

v2 = -0.203 m/s

A)

the magnitude of final velocity of 0.151 kg is 3.3 m/s

B)

magnitude of velocity of 0.302 kg is 0.203 m/s